0.45g of a dibasic acid is completely neutralized with 100ml N/10 NaOH. The molecular weight of acid is
1
Expert's answer
2020-10-19T14:28:40-0400

Solution.

"H2X + 2NaOH = 2NaX + 2H2O"

"N(NaOH) = \frac{1}{10} = 0.1 N"

"n(NaOH) = 0.1 \times 0.1 = 0.01 mol"

"n(H2X) = \frac{n(NaOH)}{2}"

n(H2X) = 0.005 mol

"M(H2X) = \frac{m(H2X)}{n(H2X)} = 90 g/mol"

Answer:

90 g/mol